Solve the system: x + y + z = 3, 2x – y + 3z = 3, x - 2y + 3z = 0
Answer by Boreal(15235)

x + y + z = 3, 2x – y + 3z = 3, x - 2y + 3z = 0
x+y+z=3
2x-y+3z=3
add
3x+4z=6
-------------
2x+2y+2z=6, double first
x-2y+3z=0
3x+5z=6
subtract
-z=0
z=0
x=2
y must be 1
(2, 1, 0)
Can be done with matrices, but eliminating one variable twice makes it two simultaneous equations and can be done that way, too.